# HybridNorm: Towards Stable and Efficient Transformer Training via Hybrid Normalization

### v1

@@ Abstract
Transformers have become the de facto architecture for a wide range of machine learning tasks, particularly in large language models (LLMs). Despite their remarkable performance, challenges remain in training deep transformer networks, especially regarding the location of layer normalization. While Pre-Norm structures facilitate easier training due to their more prominent identity path, they often yield suboptimal performance compared to Post-Norm. In this paper, we propose HybridNorm, a straightforward yet effective hybrid normalization strategy that integrates the advantages of both Pre-Norm and Post-Norm approaches. Specifically, HybridNorm employs QKV normalization within the attention mechanism and Post-Norm in the feed-forward network (FFN) of each transformer block. This design not only stabilizes training but also enhances performance, particularly in the context of LLMs. Comprehensive experiments in both dense and sparse architectures show that HybridNorm consistently outperforms both Pre-Norm and Post-Norm approaches, achieving state-of-the-art results across various benchmarks. These findings highlight the potential of HybridNorm as a more stable and effective technique for improving the training and performance of deep transformer models. Code is available at https://github.com/BryceZhuo/HybridNorm.
Machine Learning, ICML
Figure 1: Training dynamics for 1.2B dense models with Pre-Norm, HybridNorm and HybridNorm under 1T training tokens. We present the training loss, validation loss, and downstream performance on HellaSwag and ARC-Easy (0.8 EMA smoothed), demonstrating that HybridNorm achieves superior performance.
@@
1 Introduction
Transformers have become the backbone of large language models (LLMs) and a wide range of machine learning applications. These architectures are capable of modeling long-range dependencies through self-attention mechanisms, which have made them the preferred choice for a variety of tasks, including language modeling, machine translation, and image processing (Devlin et al., 2019; Brown et al., 2020; Dosovitskiy et al., 2021). However, as transformer models become deeper and more complex, ensuring stable training remains a significant challenge. One critical factor that influences training stability is the choice of normalization methods, which is crucial for mitigating issues such as internal covariate shift and gradient instability (Xiong et al., 2020). Effectively addressing these challenges is crucial for fully harnessing the potential of deep transformer models in large-scale applications.
In transformers, Layer Normalization (LayerNorm) (Ba et al., 2016) plays a central role in stabilizing training by normalizing the activations within each layer. The two predominant strategies for applying LayerNorm are Pre-Layer Normalization (Pre-Norm) and Post-Layer Normalization (Post-Norm), each with its respective benefits and trade-offs. In the Pre-Norm architecture, normalization is applied before the residual addition, resulting in a more prominent identity path that facilitates faster convergence and more stable gradients (Xiong et al., 2020). This design is particularly advantageous when training deep models, as it helps mitigate gradient-related issues that can arise during backpropagation. However, while Pre-Norm can stabilize training, it often leads to inferior final performance compared to Post-Norm (Xie et al., 2023). In contrast, Post-Norm applies normalization after the residual connection, resulting in stronger regularization effects, which contribute to improved model performance. This approach has been shown to improve the generalization ability of transformers, particularly in very deep networks (Wang et al., 2024).
Despite the benefits of each approach, there is an inherent trade-off between training stability and final model performance. Pre-Norm structures typically stabilize training but may underperform in terms of generalization, while Post-Norm architectures provide better performance but can be more difficult to train, especially in deep models. To reconcile these trade-offs, we propose a hybrid normalization method that applies QKV normalization in the attention mechanism and Post-Norm in the feed-forward network (FFN), which is named as HybridNorm. The QKV normalization in the attention mechanism stabilizes the flow of information between layers by normalizing the query, key, and value components, while Post-Norm in the FFN ensures effective depth in the deeper layers of the transformer.
Through extensive experiments on large-scale models, we validate the effectiveness of our approach. Our results show that the hybrid normalization method significantly outperforms both Pre-Norm and Post-Norm across multiple benchmarks, providing a stable training process and improved model performance. Specifically, our method achieves superior results in the context of LLMs, where the benefits of both normalization schemes are most pronounced. We believe that this hybrid approach offers a promising solution for enhancing the training stability and performance of deep transformer architectures, particularly in the rapidly evolving field of LLMs.
The main contributions of this paper can be summarized as follows.
  • We provide an empirical analysis of the proposed hybrid normalization method, demonstrating its potential benefits in terms of gradient flow stability, regularization effects, and model robustness. This analysis highlights how our approach addresses the key challenges posed by deep transformer architectures.
  • Through extensive experiments on large-scale models, we empirically validate the effectiveness of our approach. Our results show that hybrid normalization significantly outperforms both Pre-Norm and Post-Norm across a variety of tasks, leading to more stable training and improved model performance, particularly in the context of large language models (LLMs).
The remainder of this paper is organized as follows. Section 2 reviews related work on normalization techniques in transformers. Section 3 introduces our proposed HybridNorm method, detailing its design and theoretical motivation. Section 4 presents the experimental setup and evaluation results, demonstrating the effectiveness of HybridNorm across various large-scale transformer models. Section 5 concludes the paper.
2 Related Work
Architecture Modifications in Transformers. Recent efforts in transformer architecture modifications have sought to optimize both the computational efficiency and the expressiveness of the model. These efforts include changes to the attention mechanism and feed-forward networks all aimed at improving performance on a variety of tasks, ranging from language modeling to vision tasks (Vaswani et al., 2017; Dosovitskiy et al., 2021). For example, Multi-head Latent Attention (MLA) (DeepSeek-AI, 2024), Mixture of Experts (MoE) (Yuksel et al., 2012). While these modifications contribute to more efficient training, they also require careful integration with other components, such as normalization layers, to maintain model stability and performance.
Normalization Types in Transformers. Normalization layers are integral to the success of deep learning models, and transformers are no exception. The most commonly used normalization technique in transformers is LayerNorm (Ba et al., 2016), which normalizes the activations of each layer independently. However, alternative methods such as RMSNorm (Zhang & Sennrich, 2019), which normalizes using root mean square statistics, have been proposed as more effective alternatives in certain settings. These methods are designed to mitigate the challenges of internal covariate shift and gradient instability, which are critical for the success of large-scale transformer models.
Normalization Settings in Attention. For training stability, QK-Norm (Henry et al., 2020; Dehghani et al., 2023) modifies the standard attention mechanism by applying normalization directly to the query (Q) and key (K) components during attention computation. Building upon this, QKV-Norm (Menary et al., 2024; Rybakov et al., 2024) extends the approach by normalizing the Query (Q), Key (K), and Value (V) components. This comprehensive normalization ensures that all critical components of the attention mechanism are normalized, resulting in enhanced stability and improved performance.
Location of Normalization Layers. Another line of research focuses on the location of the normalization. The choice between Pre-Norm and Post-Norm architectures has been widely studied in the transformer literature (Vaswani et al., 2017; Klein et al., 2017; Wang et al., 2019). Pre-Norm, where normalization is applied before the residual connection, has been shown to be more stable in deep networks and accelerates convergence (Xiong et al., 2020). Although Post-Norm is more challenging to train, it tends to deliver better final performance by normalizing after the residual connection (Liu et al., 2020). DeepNorm (Wang et al., 2024) was proposed as a strategy to address training instability in deep transformers, which scales the residual connections by a carefully chosen factor to improve gradient flow and mitigate exploding or vanishing gradients. The method most similar to ours is Mix-LN (Li et al., 2025), which applies Post-Norm to the earlier layers and Pre-Norm to the deeper layers, achieving improved training stability and better performance. In contrast, our HybridNorm integrates Pre-Norm and Post-Norm within each transformer block.
3 Method
In this section, we first review the two predominant normalization strategies in transformer architectures: Post-Norm and Pre-Norm. We then introduce our proposed hybrid normalization method, HybridNorm, and provide its formal definition.
3.1 Preliminaries
@@
Scaled Dot-Product Attention. The scaled dot-product attention computes the attention scores between the Query (Q) and Key (K) matrices, scaled by the square root of the key dimension $d_{k}$, and applies these scores to the Value (V) matrix. The formulation is expressed as
@@
Multi-Head Attention. Multi-head attention (MHA) extends the scaled dot-product attention mechanism by splitting the query, key, and value matrices into $h$ heads, each of size $d_{k}=d/h$. Each head independently computes attention scores, and the outputs are concatenated and linearly projected to the original dimension,
@@
where $\mathrm{head}_{i}=\mathrm{attn}(Q_{i},K_{i},V_{i})$ for $i=1,2,\dots,h$, $\{\bullet_{i}\}_{i=1}^{h}=\mathrm{Split}(XW_{\bullet})$ for $\bullet\in\{Q,K,V\}$, and $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ are learnable parameters. By enabling the model to focus on different subspaces of the input representation, MHA enhances the transformer’s capacity to capture diverse patterns in the input sequence.
Feed-Forward Networks (FFN). In transformers, the Feed-Forward Network (FFN) is applied independently to each position of the input sequence.
@@ 3.2 Post-Norm and Pre-Norm
Post-Norm. Post-Norm applies the normalization layer after the residual connection in each transformer sub-layer. Formally, the output of Post-Norm can be expressed as
$Y^{l}$ $=\mathrm{Norm}(\mathrm{MHA}(X^{l})+X^{l}),$ (3)
$X^{l+1}$ $=\mathrm{Norm}(\mathrm{FFN}(Y^{l})+Y^{l}),$ (4)
where $\mathrm{Norm}$ denotes RMSNorm (Zhang & Sennrich, 2019) or LayerNorm (Ba et al., 2016).
Pre-Norm. In contrast, Pre-Norm normalizes the input to the sub-layer, which allows for a more prominent identity path. The output of Pre-Norm is given by
$Y^{l}$ $=\mathrm{MHA}(\mathrm{Norm}(X^{l}))+X^{l},$ (5)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (6)
This structure facilitates better gradient flow and stable convergence, particularly for deep models. However, its reliance on normalization before the residual connection can lead to suboptimal performance compared to Post-Norm, as the normalization does not account for the interaction between the residual connection and the sub-layers output.
3.3 HybridNorm
To address the trade-offs between Post-Norm and Pre-Norm, we propose HybridNorm, a hybrid normalization strategy that integrates their strengths. Specifically, HybridNorm combines QKV-Norm (Menary et al., 2024; Rybakov et al., 2024) in MHA and Post-Norm in FFN.
QKV Normalization in Attention. In the attention mechanism, the query, key, and value matrices are normalized individually before computing the attention output. The normalized QKV matrices are then used in the scaled dot-product attention. QKV-Norm enhances the stability of model training and leads to improved downstream performance. Formally, attention with QKV-Norm is defined as
$\mathrm{attn}_{QKV}(Q,K,V)$ (7)
$=\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(V).$
@@ 3.3 HybridNorm
HybridNorm Architecture. Combining the above, the overall output of a transformer block with HybridNorm can be expressed as
$Y^{l}$ $=\mathrm{MHA}_{QKV}(X^{l})+X^{l},$ (8)
$X^{l+1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (9)
The architecture illustration can be found in Figure 2(d) and the pseudocode is shown in Algorithm 1. By integrating QKV normalization in the attention mechanism and Post-Norm in the FFN, HybridNorm achieves stable training dynamics and enhanced final performance. The theoretical gradient analysis can be found in Appendix A.
Remark 3.1.
The method most closely related to ours is Mix-LN (Li et al., 2025), which applies Post-Norm to the earlier layers and Pre-Norm to the deeper layers, resulting in enhanced training stability and performance. In contrast, our proposed HybridNorm integrates Pre-Norm and Post-Norm within each transformer block, providing a unified approach to leverage the benefits of both normalization strategies. Moreover, experiments demonstrate that HybridNorm achieves superior downstream performance compared to Mix-LN (see Table 4).
Special Treatment of First Block. Inspired by prior work (DeepSeek-AI, 2024), which employs the Mixture of Experts (MoE) architecture with specialized handling of the first layer, we explore the impact of introducing specialized normalization to the first transformer block. In our approach, the first layer of the transformer is treated differently by applying Pre-Norm on MHA and FFN, while maintaining QKV-Norm. Specifically, the structure of our first layer is defined as
$Y^{0}$ $=\mathrm{MHA}_{QKV}(\mathrm{Norm}(X^{0}))+X^{0},$ (10)
$X^{1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$ (11)

### v3

@@ Abstract
Transformers have become the de facto architecture for a wide range of machine learning tasks, particularly in large language models (LLMs). Despite their remarkable performance, challenges remain in training deep transformer networks, especially regarding the position of layer normalization. While Pre-Norm structures facilitate more stable training owing to their stronger identity path, they often lead to suboptimal performance compared to Post-Norm. In this paper, we propose HybridNorm, a simple yet effective hybrid normalization strategy that integrates the advantages of both Pre-Norm and Post-Norm. Specifically, HybridNorm employs QKV normalization within the attention mechanism and Post-Norm in the feed-forward network (FFN) of each transformer block. We provide both theoretical insights and empirical evidence demonstrating that HybridNorm improves gradient flow and model robustness. Extensive experiments on large-scale transformer models, including both dense and sparse variants, show that HybridNorm consistently outperforms both Pre-Norm and Post-Norm approaches across multiple benchmarks. These findings highlight the potential of HybridNorm as a more stable and effective technique for improving the training and performance of deep transformer models.
[…]
1 Introduction
Transformers have become the backbone of large language models (LLMs) and a wide range of machine learning applications. These architectures are capable of modeling long-range dependencies through self-attention mechanisms, which have made them the preferred choice for a variety of tasks, including language modeling, machine translation, and image processing [8, 3, 10]. However, as transformer models become deeper and more complex, ensuring stable training remains a significant challenge. One critical factor that influences training stability is the choice of normalization methods, which is crucial for mitigating issues such as internal covariate shift and gradient instability [43]. Effectively addressing these challenges is crucial for fully harnessing the potential of deep transformer models in large-scale applications.
In transformers, Layer Normalization (LayerNorm) [1] plays a central role in stabilizing training by normalizing the activations within each layer. The two predominant strategies for applying LayerNorm are Pre-Layer Normalization (Pre-Norm) and Post-Layer Normalization (Post-Norm), each with its respective benefits and trade-offs. In the Pre-Norm architecture, normalization is applied before the residual addition, resulting in a more prominent identity path that facilitates faster convergence and more stable gradients [43]. This design is particularly advantageous when training deep models, as it helps mitigate gradient-related issues that can arise during backpropagation. However, while Pre-Norm can stabilize training, it often leads to inferior final performance compared to Post-Norm [42]. In contrast, Post-Norm applies normalization after the residual connection, resulting in stronger regularization effects, which contribute to improved model performance. This approach has been shown to improve the generalization ability of transformers, particularly in very deep networks [39]. Further discussion of related work is provided in Appendix 7..
Despite the benefits of each approach, there is an inherent trade-off between training stability and final model performance. Pre-Norm structures typically stabilize training but may underperform in terms of generalization, while Post-Norm architectures provide better performance but can be more difficult to train, especially in deep models. To reconcile these trade-offs, we propose a hybrid normalization method that applies QKV normalization in the attention mechanism and Post-Norm in the feed-forward network (FFN), which is named as HybridNorm. The QKV normalization in the attention mechanism stabilizes the flow of information between layers by normalizing the query, key, and value components, while Post-Norm in the FFN ensures the effective depth of the transformer.
Through extensive experiments on large-scale models, we validate the effectiveness of our approach. Our results show that the hybrid normalization method significantly outperforms both Pre-Norm and Post-Norm across multiple benchmarks, providing a stable training process and improved model performance. We believe that this hybrid approach offers a promising solution for enhancing the training stability and performance of deep transformer architectures, particularly in the rapidly evolving field of LLMs. The main contributions of this paper can be summarized as follows:
  • We present both theoretical and empirical analyses of HybridNorm, demonstrating its advantages in enhancing gradient flow stability and improving model robustness. Our findings underscore the method’s effectiveness in mitigating core challenges inherent to deep transformer architectures.
  • Through extensive experiments on large-scale models, we empirically validate the effectiveness of our approach. Our results show that hybrid normalization significantly outperforms both Pre-Norm and Post-Norm across a variety of tasks, leading to more stable training and improved model performance, particularly in the context of LLMs.
2 Preliminaries
@@
Scaled Dot-Product Attention
@@
The scaled dot-product attention computes the attention scores between the Query (Q) and Key (K) matrices, scaled by the square root of the key dimension $d_{k}$, and applies these scores to the Value (V) matrix. The formulation is expressed as
@@
Multi-Head Attention
Multi-head attention (MHA) extends the scaled dot-product attention mechanism by splitting the query, key, and value matrices into $h$ heads, each of size $d_{k}=d/h$. Each head independently computes attention scores, and the outputs are concatenated and linearly projected to the original dimension,
@@
where $\mathrm{head}_{i}=\mathrm{attn}(Q_{i},K_{i},V_{i})$ for $i=1,2,\dots,h$, $\{\bullet_{i}\}_{i=1}^{h}=\mathrm{Split}(XW_{\bullet})$ for $\bullet\in\{Q,K,V\}$, and learnable parameters $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ By enabling the model to focus on different subspaces of the input representation, MHA enhances the transformer’s capacity to capture diverse patterns in the input sequence.
@@ 2.1 Post-Norm and Pre-Norm
Post-Norm
Post-Norm applies the normalization layer after the residual connection in each transformer sub-layer. Formally, the output of Post-Norm can be expressed as
$Y^{l}$ $=\mathrm{Norm}(\mathrm{MHA}(X^{l})+X^{l}),\quad X^{l+1}=\mathrm{Norm}(\mathrm{FFN}(Y^{l})+Y^{l}),$ (3)
where $\mathrm{Norm}$ denotes RMSNorm [47] or LayerNorm [1].
Pre-Norm
In contrast, Pre-Norm normalizes the input to the sub-layer, which allows for a more prominent identity path. The output of Pre-Norm is given by
$Y^{l}$ $=\mathrm{MHA}(\mathrm{Norm}(X^{l}))+X^{l},\quad X^{l+1}=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+Y^{l}.$ (4)
This structure facilitates better gradient flow and stable convergence, particularly for deep models. However, its reliance on normalization before the residual connection can lead to suboptimal performance compared to Post-Norm, as the normalization does not account for the interaction between the residual connection and the sub-layers output. An analysis of the fundamental differences between the two approaches is provided in the Appendix 11.
3 HybridNorm
To address the trade-offs between Post-Norm and Pre-Norm, we propose HybridNorm, a hybrid normalization strategy that integrates their strengths. Specifically, HybridNorm combines QKV-Norm [22, 30] in MHA and Post-Norm in FFN.
QKV Normalization in Attention
In the attention mechanism, the query, key, and value matrices are normalized individually before computing the attention output. The normalized QKV matrices are then used in the scaled dot-product attention. QKV-Norm enhances the stability of model training and leads to improved downstream performance. Formally, attention with QKV-Norm is defined as
$\mathrm{attn}_{QKV}(Q,K,V)=\mathrm{softmax}\left(\frac{\mathrm{Norm}(Q)\mathrm{Norm}(K)^{\top}}{\sqrt{d_{k}}}\right)\mathrm{Norm}(V).$ (5)
@@ 3 HybridNorm
HybridNorm Architecture
Combining the above, the overall output of a transformer block with HybridNorm can be expressed as
$Y^{l}$ $=\mathrm{MHA}_{QKV}(X^{l})+X^{l},\quad X^{l+1}=\mathrm{FFN}(\mathrm{Norm}(Y^{l}))+\mathrm{Norm}(Y^{l}).$ (6)
The architecture illustration can be found in Figure 1(d) and the pseudocode is shown in Algorithm 1. By integrating QKV normalization in the attention mechanism and Post-Norm in the FFN, HybridNorm achieves stable training dynamics and enhanced final performance. The theoretical gradient analysis can be found in Appendix 8.
Remark 1.
The method most closely related to ours is Mix-LN [20], which applies Post-Norm to the earlier layers and Pre-Norm to the deeper layers, resulting in enhanced training stability and performance. In contrast, our proposed HybridNorm integrates Pre-Norm and Post-Norm within each transformer layer, thereby providing a more uniform approach across different layers to leverage the benefits of both normalization strategies. Moreover, experiments demonstrate that HybridNorm achieves superior downstream performance compared to Mix-LN (see Table 4 & Table 8).
Special Treatment of First Block
Inspired by prior work [6], which employs the Mixture of Experts (MoE) architecture with specialized handling of the first layer, we explore the impact of introducing specialized normalization to the first transformer block. In our approach, the first layer of the transformer is treated differently by applying Pre-Norm on MHA and FFN, while maintaining QKV-Norm. Specifically, the structure of our first layer is defined as
$Y^{0}$ $=\mathrm{MHA}_{QKV}(\mathrm{Norm}(X^{0}))+X^{0},\quad X^{1}=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$ (7)
@@
(a) Layer gradient norm at step 1.
(b) Layer gradient norm at step 100.
Figure 2: Gradient norm of Pre-Norm, Post-Norm, and HybridNorm at different training steps.
4 Theoretical Analysis
4.1 Benefits of Hybrid Method
To gain deeper insights into the stability introduced by HybridNorm, we follow the approach of [35, 20] and analyze the evolution of gradient norms throughout training iterations. Suppose $x$ and $F$ are the input and the sublayer of Transformer, respectively. The output of Post-Norm is $y_{Post}=\mathrm{Norm}(x+F(x))$ and the output of Pre-Norm is $y_{Pre}=x+F(\mathrm{Norm}(x))$. Then we have
$\frac{\partial y_{Post}}{\partial x}$ $=\frac{\partial\mathrm{Norm}(x+F(x))}{\partial(x+F(x))}\left(I+\frac{\partial F(x)}{\partial x}\right),$ (8)
$\frac{\partial y_{Pre}}{\partial x}$ $=I+\frac{\partial F(\mathrm{Norm}(x))}{\partial\mathrm{Norm}(x)}\frac{\partial\mathrm{Norm}(x)}{\partial x},$ (9)
where the gradient of normalization is $\frac{\partial\mathrm{Norm}(x)}{\partial x}=\alpha\odot\frac{\sqrt{d}}{||x||_{2}}\left(I-\frac{xx^{\top}}{||x||_{2}^{2}}\right)$. The gradient of Post-Norm is the product of two gradients, one of which is the normalization. If the spectral radius of the normalization gradient is less than 1, it causes gradient vanishing. In Pre-Norm, the residual connection is isolated from the normalization, ensuring that gradients retain a lower bound, thereby preventing gradient vanishing. To ensure relatively stable gradients, a natural idea is to use both types of normalization within a Transformer block, leading to Pre-Post (Pre-Norm in MHA and Post-Norm in FFN) and Post-Pre. From Table 4, we find that Pre-Post achieves the best performance, which is why we adopt this. And placing Pre-Norm in MHA with QKV-Norm further enhances performance.
In Figure 2, we compare the gradient norms of Pre-Norm, Post-Norm, and HybridNorm at steps 1 and 100. The results indicate that Pre-Norm tends to exhibit gradient explosion in deeper models, while Post-Norm suffers from vanishing gradients, both of which hinder effective optimization. In contrast, HybridNorm maintains a well-balanced gradient flow throughout training, effectively mitigating these issues. An intuitive understanding is that Pre-Norm tends to amplify gradients, while Post-Norm diminishes them. HybridNorm alternates between these two normalization strategies, leading to more stable gradient propagation during backpropagation and effectively preventing gradient explosion or vanishing. This balanced gradient propagation contributes to smoother optimization dynamics and faster convergence, further reinforcing the effectiveness of HybridNorm in stabilizing training.
4.2 Benefits of QKV-Norm
Theoretically, we study how QKV-Norm affects the gradient flow during backpropagation, which is crucial for training stability in deep transformer models. Our analysis reveals that QKV-Norm helps decouple gradients between different weight matrices, thereby stabilizing training.
Theorem 1 (Informal version of Theorem 2).
Suppose the the output of the attention is $S$, the input $X\in\mathbb{R}^{s\times d}$, parameters $W_{Q},W_{K},W_{V},W_{O}^{\top}\in\mathbb{R}^{d\times d_{k}}$. For the attention with Pre-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(\|W_{V}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{F}\right),$
$\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}$ $=\mathcal{O}\left(\|W_{K}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}=\mathcal{O}\left(\|W_{Q}\|_{2}\|W_{V}\|_{2}\|W_{O}\|_{2}\right).$
For the attention with Pre-Norm and QK-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(\|W_{V}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{F}\right),\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}=\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}=\mathcal{O}\left(\|W_{V}\|_{2}\|W_{O}\|_{2}\right).$
For the attention with QKV-Norm, we have
$\left\|\frac{\partial S}{\partial W_{O}}\right\|_{F}$ $=\mathcal{O}\left(1\right),\left\|\frac{\partial S}{\partial W_{V}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{2}\right),\left\|\frac{\partial S}{\partial W_{Q}}\right\|_{F}=\left\|\frac{\partial S}{\partial W_{K}}\right\|_{F}=\mathcal{O}\left(\|W_{O}\|_{2}\right).$
The above theorem is an informal version of Theorem 2, with a more precise statement and proof provided in Appendix 8. In attention with Pre-Norm, gradients of weights exhibit strong dependencies on other weights; for instance, $W_{Q}$ and $W_{K}$ are influenced by all three other weights but not by themselves. In contrast, with QKV-Norm, the gradient of each weight depends at most on itself and $W_{O}$. This indicates that the gradient in Pre-Norm is more tightly coupled with other weights compared to QKV-Norm, while Pre-Norm with QK-Norm lies between the two. Consequently, during training, if the norm of a certain weight becomes excessively large, it is harder to control in Pre-Norm, leading to increased gradient magnitude. This creates a vicious cycle that may cause model collapse. In contrast, QKV-Norm alleviates this issue and significantly improves training stability. In summary, the degree of gradient coupling follows: Pre-Norm $>$ Pre-Norm with QK-Norm $>$ QKV-Norm; whereas training stability follows the reverse: Pre-Norm $<$ Pre-Norm with QK-Norm $<$ QKV-Norm.